# Local Thermal Operations and Classical Communication

### v2

@@ 4 Framework of Local Thermal Operations and Classical Communication
4.1 Local thermal operations and preservation of product Gibbs state
@@
4.2 Classical communication and Landauer cost of measurement
The final ingredient one requires before going into the framework of LTOCC is the problem of local measurements and classical communication between the parties, as there are certain subtleties that need to be taken into account, especially the thermodynamic cost of measuring the system in the energy eigenbasis. Such measurements can be emulated using the SLTO framework given memory erasure after each round of measurement and forwarding of information, which we demonstrate explicitly in D. This result is similar in spirit to the incoherently conditioned thermal operations as a subset of TO, as demonstrated in [58]. Below, we lay down more heuristic arguments and point to a possible hidden cost connected with Landauer erasure.
The energy-incoherent quantum states $\rho=\sum_{i}p_{i}\;\outerproduct{E_{i}}{E_{i}}$ | can be interpreted as a probabilistic combination or a statistical ensemble of pure Hamiltonian eigenstates. The projective measurement onto the energy eigenbasis collapses it to a random eigenstate |$\ket{E_{i}}$ with probability $𝐸𝑖p_{i}=\expectationvalue{\rho}{E_{i}}$, thus single-shot measurement is not a thermal operation – it can project a Gibbs state onto a high or low energy eigenstate, resulting in a gain of free energy, otherwise inaccessible by thermal operations alone. However, statistically, incoherent measurements leave diagonal states invariant $\sum_{i}|E_{i}\rangle\langle E_{i}|\langle E_{i}|\rho|E_{i}\rangle=\sum_{i}p_{i}\;\outerproduct{E_{i}}{E_{i}}=\rho$. Thus while performing measurement on consecutive states from a Gibbs ensemble, with respect to energy eigenbasis $N$ times, the discrepancy between an ensemble of measurement outcomes and initial distribution vanishes on average as we go into the limit of a large number of rounds, $N\rightarrow\infty$.
@@ 4.2 Classical communication and Landauer cost of measurement
Additionally, we note that arguments made by Watanabe and Takagi in [58] do not translate to our setting, as Alice and Bob do not have access to shared bath, and both systems may be in different temperatures. Thus, in order to formalise heuristic considerations put forward above we explicitly consider measurement in energy eigenbasis as a channel $\Phi:\mathcal{H}_{A}\mapsto\mathcal{H}_{A}\otimes\mathcal{M}$ represented by a set of operators $K_{i}=\quantity(\ket{E_{i}}\otimes\ket{i}_{\mathcal{M}})\bra{E_{i}}$ | for all $i$, which imprint the detected information about the energy level $E_{i}$ on an auxiliary (classical) memory system $\mathcal{M}$ and assume that energies are non-degenerate. It is straightforward to see that after discarding memory the map returns pinched state $𝐸𝑖\text{Tr}_{\mathcal{M}}\quantity[\Phi(\rho)]=\mathcal{P}(\rho)=\sum_{i}\expectationvalue{\rho}{E_{i}}\outerproduct{E_{i}}{E_{i}}$ |. Pinching map, in general, preserves all energy-incoherent states and, in particular, the Gibbs state. Furthermore, the use of classical information obtained from measurement to apply a classically-conditioned map is described by channel $\Xi:\mathcal{M}\otimes\mathcal{H}_{B}\mapsto\mathcal{M}\otimes\mathcal{H}_{B}$ defined as $\Xi(\mu\otimes\sigma)=\sum_{i}\langle i|\mu|i\rangle|i\rangle\langle i|\otimes\Xi_{i}(\sigma)$, with $\quantity{\Xi_{i}:\mathcal{H}_{B}\mapsto\mathcal{H}_{B}}$ being a collection of CPTP maps. From this discussion one can infer that any further manipulation of subsystems $\mathcal{M}$ and $\mathcal{H}_{B}$ cannot alter the marginal belonging to the system $\mathcal{H}_{A}$:
$\text{Tr}_{B,\mathcal{M}}\quantity[(\mathbb{I}_{A}\otimes\Xi)\,\circ\,(\Phi\otimes\mathbb{I}_{B})\quantity(\rho\otimes\sigma)]=\text{Tr}_{\mathcal{M}}\quantity[\Phi(\rho)]=\mathcal{P}(\rho)=\rho.$
@@
Furthermore, let us highlight that one could extend the possible measurements to more general POVMs, e.g. with operators of the form $K_{i}=\sum_{j}a_{ij}\outerproduct{E_{j}}{E_{j}}$ | such that for all $j$ $\sum_{i}a_{ij}^{2}=1$, while satisfying the Gibbs preservation condition. There are three points to be made in this context. Firstly, when restricted to energy-incoherent states, such operations can be easily achieved equivalently by classical post-processing on the memory side. The second, more important aspect is that such measurements would in general provide partial coherence and/or entanglement preservation while giving only imperfect information about the system. Thirdly, althou such energy incoherent measurements preserve Gibbs state, in case of general quantum states it may not be strong enought restriction to consider it “thermal”, due to the distinction between thermal operations and Gibbs-preserving operations, which emerge on the quantum. However, due to their connection to coherence such measurements fall outside of the scope of this manuscript, and will be considered in detail in a future work [61].
Finally, because statistically neither (conditional) thermal operations nor measurement in energy eigenbasis can create any coherences, the description of energy incoherent states by probability distributions is complete, and instead of quantum channels, we can continue our discussion using stochastic maps.
4.3 Framework outline
@@
[…]
$M_{ij,kl}=T_{ikl}^{(A)}\Lambda_{jl}^{(B)}\leavevmode\nobreak$ (17)
@@
$\forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}^{(A)}=\gamma_{i}^{(A)}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{k}\sum_{l}T_{jkl}^{(B)}\gamma_{l}^{(B)}=\gamma_{j}^{(B)}\;.$ (18)
@@
$M=\prod_{a}M^{(\alpha)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ \leavevmode\nobreak\ M_{ij,kl}^{(\alpha)}=\Lambda_{ik}^{(A,\alpha)}T_{jkl}^{(B,\alpha)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ikl}^{(A,\alpha)}\Lambda_{jl}^{(B,\alpha)}\leavevmode\nobreak\ ,$ (19)
with $\alpha$ indexing the consecutive rounds. We denote all operations realizable using $n$-round protocol of the form (19) as $\text{LTOCC}_{n}$. A schematic representation of this multiround protocol is provided in Fig. 3(a).
@@
$r_{ij,m_{1}\ldots m_{\alpha}}^{\alpha}=\left[\mathbf{M^{(\alpha)}}(\mathbf{r}^{(\alpha-1)})\right]_{ijm_{1}\ldots m_{\alpha}}=\sum_{kl}{T^{(A,\alpha)}}_{i}^{k;\,m_{1}\ldots m_{\alpha-1}}\delta_{k,m_{\alpha}}\;{T^{(B,\alpha)}}_{j}^{l;\,m_{1}\ldots m_{\alpha-1}m_{\alpha}}\;r_{kl,m_{1}\ldots m_{\alpha-1}}^{(\alpha-1)}$ (21)
where $m_{i},\ldots,m_{\alpha}$ are indices related to the memory registers, $k,l$ are system-related input state indices, and $i,j$ are system-related output-state indices. Here, $\delta_{k,m_{\alpha}}$ is a measurement on Alice’s subsystem, recorded on a new classical register $m_{\alpha}$. This is followed by thermal operation ${T^{(A)}}_{i}^{k;\,m_{1}\ldots m_{\alpha-1}}$ conditioned on all previous measurement results $m_{1}\ldots,m_{\alpha-1}$. Furthermore ${T^{(B)}}_{j}^{l;\,m_{1}\ldots m_{\alpha-1}m_{\alpha}}$ is a thermal operation on Bob’s side conditioned by the measurement’s outcome in previous rounds $m_{1},\ldots,m_{\alpha-1}$ and new measurement result on the Alice side $m_{\alpha}$. We provide a diagrammatic representation of this protocol in Fig.  3(b). Additional detailed discussion of 1- and 2-round LTOCC protocols with and without protocols is provided in C.
(a) $\text{LTOCC}_{2}$
@@
@C=1.0em @R=1.0em @!R
[…]
(b) $\text{LTOCC}_{2}+\text{M}$
[…]
@C=1.0em @R=1.0em @!R
[…]
@@
In a slight abuse of terminology, we refer to this particular restricted scenario as 1-round parallel LTOCC, as it does not involve any feedback. This procedure is depicted in Fig.  4.
[…]
@C=1.0em @R=1.0em @!R
[…]
@@
$M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}\Lambda_{ik}^{(A,\alpha)}T_{jkl}^{(B,\alpha)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}T_{ikl}^{(A,\alpha)}\Lambda_{jl}^{(B,\alpha)}$ (23)
@@
[…]
@@
Proof of this theorem is presented in D. ∎
We stress that the following theorem describes only protocols without memory. If we allow the use of outcomes of some measurements in multiple rounds, the class of obtained protocols might be noticeably more powerful, as presented in the C.
@@
$M(\gamma^{(A)}\otimes\gamma^{(B)})=\gamma^{(A)}\otimes\gamma^{(B)}$
We stress that the above corollary describes the implication in only one direction. The implication in the opposite direction would imply that LTOCC $=$ SLTO.
@@
$\forall_{k}\sum_{l}T_{ikl}^{(A)}\gamma_{l}=\gamma_{i}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}=\gamma_{j}$ (25)
@@
Should the above conjecture, or its stronger version without shared randomness or extension to genuinely quantum states, hold true, it would have interesting consequences, especially taking into account that considerations in [33] have been carried out appealing only energy-incoherent states – finite-size engine operating with Carnot efficiency could be realised with a (potentially finite-round) LTOCC protocol without appealing to its quantum properties. Furthermore, in LTOCCn +R all interactions between subsystems and the extension between thermal baths are purely classical, which is not the case for SLTO. Hence, the above result would indicate that at least for energy-incoherent states, the introduction of coherent interactions does not increase the set of possible operations.
@@
5 Basic building blocks – thermal and bithermal tensors
@@
In the finite temperature case, $\beta>0$, the situation is more convoluted, even when restricted to standard thermal operations [29]. As presented in [62], the structure of vertices for the thermal equivalent of Birkhoff polytope depends on the relations between system’s energies and temperature. Thus we start the discussion by presenting a simple, yet comprehensive, example of a bithermal operation on a 2-level system.
@@
$T_{ijk}^{(1)}=(T_{0,j,k}^{(1)}|T_{1,j,k}^{(1)})=\left(\begin{matrix}1-g_{1,0}&1\\ 1&0\\ \end{matrix}\right.\left|\begin{matrix}g_{1,0}&0\\ 0&1\\ \end{matrix}\right)\leavevmode\nobreak\ ,$ (27)
$T_{ijk}^{(2)}=(T_{0,j,k}^{(2)}|T_{1,j,k}^{(2)})=\left(\begin{matrix}1-g_{1,0}(1-g_{1,0})&1-g_{1,0}\\ 1-g_{1,0}&1\\ \end{matrix}\right.\left|\begin{matrix}g_{1,0}(1-g_{1,0})&g_{1,0}\\ g_{1,0}&0\\ \end{matrix}\right)\leavevmode\nobreak\ .$ (28)
@@
$T_{ijk}^{(2)}=\sum_{l}S_{il}^{(\beta)}T_{ljk}^{(1)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ilk}^{(1)}S_{lj}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ljl}^{(1)}S_{lk}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak$ (29)
where $\text{where }\leavevmode\nobreak\ S^{(\beta)}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\end{matrix}\right)\leavevmode\nobreak\ .$
@@
5.1 Extremal points of thermal “Birkhoff” polytope in infinite and large temperatures
@@
$T_{ijk}=\widetilde{T}_{ijk}+\beta A_{ijk}^{(1)}+O(\beta^{2})\leavevmode\nobreak\ .$ (30)
@@
${\gamma_{E}}_{i}=\frac{e^{-\beta E_{i}}}{\sum_{j}e^{-\beta E_{j}}}=\frac{1}{d}\left(1+\beta\frac{\sum_{m}E_{m}}{d}\right)-\beta\frac{E_{i}}{d}+O(\beta^{2})$ (31)
@@
$\sum_{i}A_{ijk}^{(1)}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{j}A_{ijk}^{(1)}=\sum_{j}\tilde{T}_{ijk}E_{j}-E_{i},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{k}A_{ijk}^{(1)}=\sum_{k}\tilde{T}_{ijk}E_{k}-E_{i}$ (32)
@@
$A_{ijk}^{(1)}\geq 0\leavevmode\nobreak\ \text{ if }\leavevmode\nobreak\ \tilde{T}_{ijk}=0.$ (33)
@@
5.2 Bithermal tensors approaching zero temperature
The opposite extreme of zero temperature $\beta\to\infty$ turns out to be surprisingly simple. As stated in [45], for temperatures low enough, if the thermal bath has any spectral gap, effectively its entire population lies in the ground state. In such a scenario, thermal operations reduce to cooling maps defined by condition $\Lambda_{ij}=0$ if $i>j$. Moreover, in the B, we present an alternative derivation of cooling maps, as a zero temperature limit $\beta\to\infty$ of thermal operations.

### v4

@@ IV Framework of Local Thermal Operations and Classical Communication
IV.1 Local thermal operations and preservation of product Gibbs state
@@
IV.2 Classical communication and Landauer cost of measurement
The final ingredient one requires before going into the framework of LTOCC is the problem of local measurements and classical communication between the parties, as there are certain subtleties that need to be discussed, especially the thermodynamic cost of measuring the system in the energy eigenbasis. Such measurements can be emulated using the SLTO framework given memory erasure after each round of measurement and forwarding of information, which we demonstrate explicitly in Appendix D. This result is similar in spirit to the incoherently conditioned thermal operations as a subset of TO, as demonstrated in [63]. Below, we lay down more heuristic arguments and point to a possible hidden cost connected with Landauer erasure.
[…]
@@ IV.2 Classical communication and Landauer cost of measurement
[…]
$\text{Tr}_{B,\mathcal{M}}\quantity[(\mathbb{I}_{A}\otimes\Xi)\,\circ\,(\Phi\otimes\mathbb{I}_{B})\quantity(\rho\otimes\sigma)]=\text{Tr}_{\mathcal{M}}\quantity[\Phi(\rho)]$
$=\mathcal{P}(\rho)=\rho.$
@@
Furthermore, let us highlight that one could extend the possible measurements to more general POVMs, e.g. with operators of the form $|K_{i}=\sum_{j}a_{ij}\outerproduct{E_{j}}{E_{j}}$ such that for all $j$ $\sum_{i}a_{ij}^{2}=1$, while satisfying the Gibbs preservation condition. There are three points to be made in this context. Firstly, when restricted to energy-incoherent states, such operations can be easily achieved equivalently by classical post-processing on the memory side. The second, more important aspect is that such measurements would in general provide partial coherence and/or entanglement preservation while giving only imperfect information about the system. Thirdly, althou such energy-incoherent measurements preserve Gibbs state, in case of general quantum states it may not be strong enought restriction to consider it “thermal”, due to the distinction between thermal operations and Gibbs-preserving operations, which emerge on the quantum level. However, due to their connection to coherence such measurements fall outside of the scope of this manuscript, and will be considered in detail in a future work [66].
Finally, because statistically neither (conditional) thermal operations nor measurement in energy eigenbasis can create any coherences, the description of energy-incoherent states by probability distributions is complete, and instead of quantum channels, we can continue our discussion using stochastic maps.
IV.3 Framework outline
@@
we obtain one-round Local Thermal Operations and Classical communication (LTOCC). [^1] Note that $\text{LTOCC}\;\text{``}\neq\text{''}\;\text{LOCC}_{d^{2}}+\text{TO}_{d^{2}}$, but rather $\text{LTOCC}\;\text{``}=\text{''}\;\text{LTO}_{d}^{\otimes 2}+\text{CC}_{d}^{2}$, with subscripts pointing to dimensions of the involved subsystems. In particular, the former would result in a theory identical to LOCC when restricted to classical states. . The mirror scenario of measurement on Bob’s side corresponds to
$M_{ij,kl}=T_{ikl}^{(A)}\Lambda_{jl}^{(B)}\penalty 10000\$(17)
@@
$\forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}^{(A)}=\gamma_{i}^{(A)}\;,\penalty 10000\ \penalty 10000\ \forall_{k}\sum_{l}T_{jkl}^{(B)}\gamma_{l}^{(B)}=\gamma_{j}^{(B)}\;.$ (18)
@@
$M=\prod_{\alpha}M^{(\alpha)}\penalty 10000\ \penalty 10000\ \text{where}\penalty 10000\ \penalty 10000\$$M_{ij,kl}^{(\alpha)}=\Lambda_{ik}^{(A,\alpha)}T_{jkl}^{(B,\alpha)}$
or $M_{ij,kl}^{(\alpha)}=T_{ikl}^{(A,\alpha)}\Lambda_{jl}^{(B,\alpha)}\penalty 10000\ ,$ (19)
with $\alpha$ indexing the consecutive rounds. We denote all operations realizable using $n$-round protocol of the form (19) as $\text{LTOCC}_{n}$. A schematic representation of this multiround protocol is provided in Fig. 3.
@@
$r_{ij,m_{1}\ldots m_{\alpha}}^{\alpha}=\left[\mathbf{M^{(\alpha)}}(\mathbf{r}^{(\alpha-1)})\right]_{ijm_{1}\ldots m_{\alpha}}$
$=$ $\sum_{kl}{T^{(A,\alpha)}}_{i}^{k;\,m_{1}\ldots m_{\alpha-1}}\delta_{k,m_{\alpha}}\;{T^{(B,\alpha)}}_{j}^{l;\,m_{1}\ldots m_{\alpha-1}m_{\alpha}}\times$
@@
where $m_{i},\ldots,m_{\alpha}$ are indices related to the memory registers, $k,l$ are system-related input state indices, and $i,j$ are system-related output-state indices. Here, $\delta_{k,m_{\alpha}}$ is a measurement on Alice’s subsystem, recorded on a new classical register $m_{\alpha}$. This is followed by thermal operation ${T^{(A)}}_{i}^{k;\,m_{1}\ldots m_{\alpha-1}}$ conditioned on all previous measurement results $m_{1}\ldots,m_{\alpha-1}$. Furthermore ${T^{(B)}}_{j}^{l;\,m_{1}\ldots m_{\alpha-1}m_{\alpha}}$ is a thermal operation on Bob’s side conditioned by the measurement’s outcome in previous rounds $m_{1},\ldots,m_{\alpha-1}$ and new measurement result on the Alice side $m_{\alpha}$. We provide a diagrammatic representation of this protocol in Fig. 3. Additional detailed discussion of 1- and 2-round LTOCC protocols with and without protocols is provided in Appendix C.
@@
Figure 3.a: $\text{LTOCC}_{2}$
Figure 3.b: $\text{LTOCC}_{2}+\text{M}$
@@
In a slight abuse of terminology, we refer to this particular restricted scenario as 1-round parallel LTOCC, as it does not involve any feedback. This procedure is depicted in Fig. 4.
@@
$M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}\Lambda_{ik}^{(A,\alpha)}T_{jkl}^{(B,\alpha)}\penalty 10000\ \penalty 10000\ \text{or}\penalty 10000\ \penalty 10000\ M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}T_{ikl}^{(A,\alpha)}\Lambda_{jl}^{(B,\alpha)}$ (23)
@@
On the other hand, the introduction of shared randomness enables one to perform arbitrary convex combinations of extremal channels $M_{ij}=\sum_{\alpha_{1},\alpha_{2}}\lambda_{\alpha_{1},\alpha_{2}}\Lambda_{ik}^{\alpha_{1}}T_{jkl}^{\alpha_{2}}$, thus enlarging the framework by enabling arbitrary convex combinations. This shows that $\text{LTOCC}+\text{R}\supset\text{LTOCC}$. ∎
@@
Proof of this theorem is given in Appendix D. ∎
We stress that the following theorem describes only protocols without memory. If we allow the use of outcomes of some measurements in multiple rounds, the class of obtained protocols might be noticeably more powerful, as presented in Appendix C.
@@
$M\quantity(\gamma^{(A)}\otimes\gamma^{(B)})=\gamma^{(A)}\otimes\gamma^{(B)}$
We stress that the above corollary describes the implication in only one direction. The implication in the opposite direction would imply that LTOCC == SLTO.
@@
$\forall_{k}\sum_{l}T_{ikl}^{(A)}\gamma_{l}=\gamma_{i}\;,\penalty 10000\ \penalty 10000\ \forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}=\gamma_{j}$ (25)
@@
Should the above conjecture, its stronger version without shared randomness or extension to genuinely quantum states, hold true, it would have interesting consequences, especially taking into account that considerations in [34] have been carried out appealing only energy-incoherent states – finite-size engine operating with Carnot efficiency could be realised with a (potentially finite-round) LTOCC protocol without appealing to its quantum properties. Furthermore, in LTOCCn +R all interactions between subsystems and by extension between thermal baths are purely classical, which is not the case for SLTO. Hence, the above result would indicate that at least for energy-incoherent states, the introduction of coherent interactions does not increase the set of possible operations. At the same time, it is worth noting that, in line with the discussion provided in Section IV.2, LTOCC protocols involve external costs connected to continued writing and erasure of classical memory used to track measurement results. Thus, the question of whether the energetic cost of memory impedes attaining Carnot limit merits further investigation.
@@
For the sake of simplicity we introduced the entire LTOCC(++M) framework in restriction to the energy-incoherent states. An illustrative example of this paradigm – implementing classical logic gates CNOT and SWAP by LTOCC is presented in Appendix F. However, they are well defined for general quantum states as well, see Appendix D. Furtheremore in Appendix E we provide partial resoults on coherences’ evolutions under LTOCC(++M).
V Basic building blocks – thermal and bithermal tensors
@@
In the finite temperature case, $\beta>0$, the situation is more convoluted, even when restricted to standard thermal operations [30]. As presented in [67], the structure of vertices for the thermal equivalent of Birkhoff polytope depends on the relations between system’s energies and temperature. Thus we start the discussion by presenting a simple, yet comprehensive, example of a bithermal operation on a two-level system.
@@
$T_{ijk}^{(1)}=(T_{0,j,k}^{(1)}|T_{1,j,k}^{(1)})=\left(\begin{matrix}1-g_{1,0}&1\\ 1&0\\ \end{matrix}\right.\left|\begin{matrix}g_{1,0}&0\\ 0&1\\ \end{matrix}\right)\penalty 10000\ ,$ (27)
$T_{ijk}^{(2)}=(T_{0,j,k}^{(2)}|T_{1,j,k}^{(2)})=\left(\begin{matrix}1-g_{1,0}(1-g_{1,0})&1-g_{1,0}\\ 1-g_{1,0}&1\\ \end{matrix}\right.\left|\begin{matrix}g_{1,0}(1-g_{1,0})&g_{1,0}\\ g_{1,0}&0\\ \end{matrix}\right)\penalty 10000\ .$ (28)
@@
$T_{ijk}^{(2)}=\sum_{l}S_{il}^{(\beta)}T_{ljk}^{(1)}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ T_{ijk}^{(2)}=\sum_{l}T_{ilk}^{(1)}S_{lj}^{(\beta)}\penalty 10000\ ,$ (29)
$T_{ijk}^{(2)}=\sum_{l}T_{ljl}^{(1)}S_{lk}^{(\beta)}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \text{where }\penalty 10000\ S^{(\beta)}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\end{matrix}\right).$
@@
V.1 Extremal points of thermal “Birkhoff” polytope in infinite and large temperatures
@@
$T_{ijk}=\widetilde{T}_{ijk}+\beta A_{ijk}^{(1)}+O(\beta^{2})\penalty 10000\ .$ (30)
@@
$\sum_{i}A_{ijk}^{(1)}=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{j}A_{ijk}^{(1)}=\sum_{j}\tilde{T}_{ijk}E_{j}-E_{i},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{k}A_{ijk}^{(1)}=\sum_{k}\tilde{T}_{ijk}E_{k}-E_{i}$ (31)
@@
$A_{ijk}^{(1)}\geq 0\penalty 10000\ \text{ if }\penalty 10000\ \tilde{T}_{ijk}=0.$ (32)
@@
V.2 Bithermal tensors approaching zero temperature
The opposite extreme of zero temperature $\beta\to\infty$ turns out to be surprisingly simple. As stated in [46], for temperatures low enough, if the thermal bath has any spectral gap, effectively its entire population lies in the ground state. In such a scenario, thermal operations reduce to cooling maps defined by condition $\Lambda_{ij}=0$ if $i>j$. Moreover, in the Appendix B, we present an alternative derivation of cooling maps, as a zero temperature limit $\beta\to\infty$ of thermal operations.